Suppose a large shipment of stereos contained 21% defectives. If a sample of size 348 is selected, what is the probability that the sample proportion will differ from the population proportion by more than 4%? Round your answer to four decimal places.
Answer #1

Solution

\sigma\hat p =  \sqrt[p ( 1 - p ) / n] = \sqrt [(0.21 * 0.79) / 348 ] = 0.0218

P(sample proportion will differ from the population proportion by more than 4%)

= 1 - P[ -0.04 / 0.0218 < ( \hat p - \mu \hat p ) / \sigma \hat p < 0.04 / 0.0218]

= 1 - P(-1.83 < z < 1.83)

= 1 - 0.9327

= 0.0673